# Robust Weight Imprinting: Insights from Neural Collapse and Proxy-Based Aggregation

### v3

@@ 1 Introduction
In machine learning applications, training models from scratch is often not viable due to limitations in data and compute. A popular solution is to apply transfer learning (transferlearning3; transferlearning2) based on foundation models (FMs) (foundationmodels) that are pre-trained on a large amount of data. To tune a FM to a novel task, e.g., for classification, a common procedure is to freeze the model parameters and replace the output layer with a new head. A particularly simple method for implementing such a new head was proposed by qi-imprinting and coined imprinting.
@@
In the original work by qi-imprinting, the last-layer weight vector of a novel class is set to the normalized average of its scaled embedding vectors, i.e., its class mean. These class means are representatives of the classes, which we denote as proxies. In general, we refer to imprinting as efficient learning methods without the need for cross-class statistics or gradient-based optimization. A plethora of studies have emerged surveying this technique by adding complexity and adaptability (fewshot-metagenweights; masked-imprinting; hypersphere-imprinting; imprint-multilabel; imprint-objdetec; multiple-projection-head-imprinting; yan2023few). Despite its many adaptations, imprinting lacks a systematic comparison that unifies them. Understanding its variations could unlock greater efficiency across many fields, making the method even more versatile and impactful.
@@
In particular, while in some practical applications, accuracy is prioritized over computational efficiency, the latter does become a critical requirement in scenarios where computational resources are severely constrained, e.g., in the chemical and polymer processing industries. Here, battery-powered edge compute is essential and frequent retraining or large-scale optimization is infeasible. Imprinting has proven particularly effective in these contexts of edge-embedded devices (hypersphere-imprinting). For instance, imprintingrobot implement a vision-based robotic force grasping with a variable-stiffness gripper that can safely handle both fragile and heavy objects by rapidly adapting to novel categories without retraining (Continual Learning (CL) setting). Industrial adoption also underscores this trend with Google’s Coral Edge TPU including an ImprintingEngine API (coralai), that allows users to add new classes from a few examples without recompiling the model. Very recently, (belal2025fsid) apply imprinting to spectrogram embeddings from IMU/EMG gait data, achieving efficient classification in low-data Human Activity Recognition (HAR) tasks. Apart from that, ncmbaseline extend imprinting to CL, establishing a simple baseline with competitive performance compared to more sophisticated state-of-the-art CL algorithms.
@@
When neural networks are trained to reach near-zero loss, their penultimate-layer embeddings collapse to the class means (papyan2020prevalence; zhu2021geometric). We investigate this phenomenon as a potential explanation for when and why imprinting works. Our analysis proves that there exists a relationship between a measure of neural collapse and the success of imprinting. Since quantification of this phenomenon is possible in a post-hoc fashion over the FM features, we believe that these insights will contribute to further development of imprinting methods, as well as, training regimes of FMs that are more suitable for transfer learning via imprinting.
@@ 2 Related Work
Weight imprinting is implemented by setting the final layer weights for the novel classes to the scaled average of the embedding vectors of their training samples and was first introduced by qi-imprinting for the few-shot learning scenario. The authors find that for up to $20$ samples, using a combination of imprinting and fine-tuning outperforms other state-of-the-art methods, including nearest neighbor algorithms. In contract, in our work we do not limit the number of samples and perform no fine-tuning on the imprinted weights to maintain efficiency.
Imprinting has been applied to object detection (imprint-objdetec; yan2023few), multi-label classification (imprint-multilabel), semantic segmentation (masked-imprinting), and combined with an attention mechanism to generate weights for the novel classes in a few-shot classification task (fewshot-metagenweights).
itsdone apply imprinting using quantile normalization to ensure statistical similarity between new and existing weights. We consider this as one normalization scheme in our framework. covidimprint apply imprinting in chest radiography for detection of COVID-19 and find that it yields better results than joint gradient descent training of all classes when only few samples are available. They speculate whether normalization is a constraint in their imprinting model.
Before the era of deep learning, first-nearestclassmean analyze the transferability of hand-crafted image features. They use a “nearest class multiple centroids” (NCMC) classifier with multiple proxies generated from a $k$-means clustering algorithm. In combination with metric learning, they compare favorably against the $m$-nearest neighbor algorithm. Our work, on the other hand, highlights efficient transfer learning provided by foundation models.
@@
Using embedding vectors from pre-trained models is a simple and widely used transfer learning approach, established in the seminal works on computer vision (donahue_decaf_2014) and natural language processing (devlin_bert_2019). kornblith_better_2019 show that pre-training performance of a model is highly correlated with the performance of the resulting embedding vectors in downstream tasks. In addition, huh_what_2016 provide insights into the required quality of pre-training data. Our work is orthogonal to these studies, since we focus on studying weight generation, normalization, and aggregation techniques applied later on for new task adaptation.
@@
Although we investigate transfer learning scenarios, we review the imprinting applications and results from CL. icarl dynamically select a subset of examples for each class and update internal representations via gradient descent. They use a nearest mean classifier (NMC) with respect to the saved examples. ncmbaseline use an NMC classifier as well and achieve good performance on CL benchmarks without any fine-tuning of the embeddings. However, they do not investigate the effect of normalization and using multiple proxies.
Findings of imprintallknn show that a simple, approximate $m$-nearest neighbor classifier outperforms existing methods in an Online CL setting when all data can be stored. In our work, however, we compare imprinting all data to a limited number of more representative proxies striving for efficiency.
@@
The phenomenon of NC was identified by papyan2020prevalence and refers to the convergence of the last-layer weight vectors to class means. It was shown that, regardless of the loss function, optimizer, batch-normalization, or regularization, NC will eventually occur (provided the training data has a balanced distribution) (zhu2021geometric; neuralcollapse2; neuralcollapse3). Nevertheless, complete neural collapse is practically unrealistic (DBLP:conf/icml/TirerHN23). In transfer learning, DBLP:conf/iclr/Galanti0H22 show that NC occurs on new samples and classes from the same distribution as the pre-training dataset, highlighting the usability of foundational models in such scenarios. In our work, we expand the survey on NC by experimenting with out-of-distribution classes belonging to different datasets and linking their degree of collapse to the success of certain imprinting strategies.
@@ 3.1 IMPRINT
We analyze the effect of weight generation (GEN), normalizations ($\texttt{NORM}=\{\texttt{NORM}\textsubscript{{pre}},\texttt{NORM}\textsubscript{{post}},\texttt{NORM}\textsubscript{{inf}}\}$), and aggregation (AGG). The IMPRINT framework depicted in fig.˜1 consists of three main building blocks: a foundation model FM, a weight generator GEN, and extendable classifier weights that are imprinted. The FM remains frozen throughout the experiments. It receives data from T as inputs and produces embedding vectors. The training process generates weight vectors for each of the $C$ classes in T. Hereby, embeddings from the FM are normalized before the generation (GEN) step according to NORMpre. The generated weight vectors per class are referred to as proxies, prototypes, or representatives (metric-proxies; prototypical; convolutional-prototypes; semanticdriftCL). These proxies are normalized according to NORMpost. As in the work by qi-imprinting, we do not use bias values. To classify the test data in T during inference, it is first embedded by the FM, normalized according to NORMinf, and finally aggregated by AGG, resulting in a predicted class label.
@@
The purpose of GEN is to determine how the embeddings of the training data in T are used to form the new weights. In contrast to qi-imprinting which only incorporates one proxy per class (the mean), we add flexibility by allowing each class to have multiple proxies as in first-nearestclassmean and enable non-linear classification. We denote the number of proxies per class as $k$, ranging between $1$ and the number of samples, and investigate the following operations conducted per class to generate those:
@@
  • k-means: $k$-means cluster centers using KMeans from sklearn (scikit-learn). $k=1$ is the same as mean.
@@
quantile normalization (quantnorm1; quantnorm2) can only be applied to generated weights. This non-linear operation distributes weights equally. Recall that if more than one class is contained in T ($c>1$), GEN is performed for each class consecutively, and the reference distribution changes accordingly. In particular, for the first class there is no reference distribution to map to. This is different from itsdone, where new weights are matched to the distribution of the original classifier weights of the FM. Since we do not consider the classes used for pre-training the FM and especially do not assume access to their last-layer weights, this is not possible in our scenario.
@@ 3.2 Quantifying Neural Collapse
Neural collapse (NC) (papyan2020prevalence) characterizes the state of the features produced by a classification neural network after training to near zero training loss. Namely, the learned embeddings of each class converge, i.e., collapse, to their class means. These globally centered class means and classifier weights form a simplex equiangular tight frame (ETF) – a collection of equal length and maximally equiangular vectors, that maximize the between-class variability. This results in an optimal linearly separable state for classification. In fig.˜3 (left), we illustrate the collapse of a FM on its pre-training data. The newly arrived data T from a different dataset is distributed more unevenly across the embedding space (right).
@@
Two important characteristics of NC are variability collapse, i.e., the within-class variability of the penultimate-layer embeddings collapses to zero, and convergence to nearest-mean-classification. We focus on variability collapse ($\mathcal{NC}_{1}$) as in zhu2021geometric. Given a foundation model FM and a finite dataset containing $C$ classes with (for simplicity) $N$ samples per class, we use its $l$-dimensional embeddings $\{\boldsymbol{\mathbf{h}}_{c,i}\}_{1\leq c\leq C,\;1\leq i\leq N}$ given by FM to define the global feature mean $\boldsymbol{\mathbf{h}}_{G}\coloneqq\frac{1}{CN}\sum_{c=1}^{C}\sum_{i=1}^{N}\boldsymbol{\mathbf{h}}_{c,i}$, class means $\overline{\boldsymbol{\mathbf{h}}}_{c}\coloneqq\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\mathbf{h}}_{c,i}$, the within-class covariance matrix $\mathbf{\Sigma}_{W}\coloneqq\frac{1}{CN}\sum_{c=1}^{C}\sum_{i=1}^{N}\bigl(\boldsymbol{\mathbf{h}}_{c,i}-\overline{\boldsymbol{\mathbf{h}}}_{c}\bigr)\bigl(\boldsymbol{\mathbf{h}}_{c,i}-\overline{\boldsymbol{\mathbf{h}}}_{c}\bigr)^{\top}$, and the between-class covariance matrix $\mathbf{\Sigma}_{B}\coloneqq\frac{1}{C}\sum_{c=1}^{C}\bigl(\overline{\boldsymbol{\mathbf{h}}}_{c}-\boldsymbol{\mathbf{h}}_{G}\bigr)\bigl(\overline{\boldsymbol{\mathbf{h}}}_{c}-\boldsymbol{\mathbf{h}}_{G}\bigr)^{\top}$ to finally compute
@@
Based on eq.˜1, an $\mathcal{NC}_{1}$ score closer to zero signifies a higher collapse. In contrast, an increase in multi-modality of data leads to a higher $\mathcal{NC}_{1}$ score (as analyzed in fig.˜11(a)). Note that this measurement is not independent of the embedding dimension $l$ and the number of classes $C$. According to NC, imprinting the mean, as originally done by qi-imprinting, is best when $\mathcal{NC}_{1}$ is small. We claim that when the data is not fully collapsed (as is often the case in practice), the scale of $\mathcal{NC}_{1}$ could guide the proxy generation method, e.g., having multiple proxies $k>1$ per class. We investigate this in section˜5.3.
@@ 3.3 Significance Testing with Critical Difference Diagrams
Our experiments compare large numbers of imprinting configurations across tasks with substantially different achievable accuracy levels. To obtain scale-independent statistical comparisons, we base our analysis on ranking (dis-)agreements rather than raw accuracy. Non-parametric rank-based tests are well suited for this setting. Following cddiag, we use the Friedman test and Wilcoxon signed-rank tests because they do not assume normality or homoscedasticity, and can be applied uniformly to any evaluation metric. Empirical evidence in cddiag further suggests that these tests have good power in classifier comparison scenarios.
Concretely, for each evaluation instance (i.e., foundation model FM and task T) we compute the classification accuracy of all configurations and assign ranks (rank 1 = highest accuracy). This yields a matrix of ranks with one row per configuration and one column per evaluation instance. On this matrix we first apply the Friedman test to assess whether there are any overall differences between configurations. If the Friedman test rejects the global null hypothesis at significance level $\alpha=0.05$, we proceed with a post-hoc, pairwise analysis. Specifically, for every pair of configurations we run a two-sided Wilcoxon signed-rank test across evaluation instances. The resulting $p$-values are corrected for multiple comparisons using the Holm step-down procedure (holm1979simple).

### v4

@@ 1 Introduction
In machine learning applications, training models from scratch is often not viable due to limitations in data and compute. A popular solution is to apply transfer learning (Bengio, 2012; Yosinski et al., 2014) based on foundation models (FMs) (Bommasani et al., 2021) that are pre-trained on a large amount of data. To tune a FM to a novel task, e.g., for classification, a common procedure is to freeze the model parameters and replace the output layer with a new head. A particularly simple method for implementing such a new head was proposed by Qi et al. (2018) and coined imprinting.
@@
In the original work by Qi et al. (2018), the last-layer weight vector of a novel class is set to the normalized average of its scaled embedding vectors, i.e., its class mean. These class means are representatives of the classes, which we denote as proxies. In general, we refer to imprinting as efficient learning methods without the need for cross-class statistics or gradient-based optimization. A plethora of studies have emerged surveying this technique by adding complexity and adaptability (Gidaris and Komodakis, 2018; Siam et al., 2019; Passalis et al., 2020; Khan et al., 2021; Li et al., 2021; Cristovao et al., 2022; Yan et al., 2023). Despite its many adaptations, imprinting lacks a systematic comparison that unifies them. Understanding its variations could unlock greater efficiency across many fields, making the method even more versatile and impactful.
@@
In particular, while in some practical applications, accuracy is prioritized over computational efficiency, the latter does become a critical requirement in scenarios where computational resources are severely constrained, e.g., in the chemical and polymer processing industries. Here, battery-powered edge compute is essential and frequent retraining or large-scale optimization is infeasible. Imprinting has proven particularly effective in these contexts of edge-embedded devices (Passalis et al., 2020). For instance, Zhu et al. (2022) implement a vision-based robotic force grasping with a variable-stiffness gripper that can safely handle both fragile and heavy objects by rapidly adapting to novel categories without retraining (Continual Learning (CL) setting). Industrial adoption also underscores this trend with Google’s Coral Edge TPU including an ImprintingEngine API (Coral, ), that allows users to add new classes from a few examples without recompiling the model. Very recently, (Belal et al., 2025) apply imprinting to spectrogram embeddings from IMU/EMG gait data, achieving efficient classification in low-data Human Activity Recognition (HAR) tasks. Apart from that, Janson et al. (2022) extend imprinting to CL, establishing a simple baseline with competitive performance compared to more sophisticated state-of-the-art CL algorithms.
@@
When neural networks are trained to reach near-zero loss, their penultimate-layer embeddings collapse to the class means (Papyan et al., 2020; Zhu et al., 2021). We investigate this phenomenon as a potential explanation for when and why imprinting works. Our analysis proves that there exists a relationship between a measure of neural collapse and the success of imprinting. Since quantification of this phenomenon is possible in a post-hoc fashion over the FM features, we believe that these insights will contribute to further development of imprinting methods, as well as, training regimes of FMs that are more suitable for transfer learning via imprinting.
@@ 2 Related Work
Weight imprinting is implemented by setting the final layer weights for the novel classes to the scaled average of the embedding vectors of their training samples and was first introduced by Qi et al. (2018) for the few-shot learning scenario. The authors find that for up to $20$ samples, using a combination of imprinting and fine-tuning outperforms other state-of-the-art methods, including nearest neighbor algorithms. In contract, in our work we do not limit the number of samples and perform no fine-tuning on the imprinted weights to maintain efficiency.
Imprinting has been applied to object detection (Li et al., 2021; Yan et al., 2023), multi-label classification (Khan et al., 2021), semantic segmentation (Siam et al., 2019), and combined with an attention mechanism to generate weights for the novel classes in a few-shot classification task (Gidaris and Komodakis, 2018).
Hosoda et al. (2024) apply imprinting using quantile normalization to ensure statistical similarity between new and existing weights. We consider this as one normalization scheme in our framework. Zhang et al. (2021) apply imprinting in chest radiography for detection of COVID-19 and find that it yields better results than joint gradient descent training of all classes when only few samples are available. They speculate whether normalization is a constraint in their imprinting model.
Before the era of deep learning, Mensink et al. (2013) analyze the transferability of hand-crafted image features. They use a “nearest class multiple centroids” (NCMC) classifier with multiple proxies generated from a $k$-means clustering algorithm. In combination with metric learning, they compare favorably against the $m$-nearest neighbor algorithm. Our work, on the other hand, highlights efficient transfer learning provided by foundation models.
@@
Using embedding vectors from pre-trained models is a simple and widely used transfer learning approach, established in the seminal works on computer vision (Donahue et al., 2014) and natural language processing (Devlin et al., 2019). Kornblith et al. (2019) show that pre-training performance of a model is highly correlated with the performance of the resulting embedding vectors in downstream tasks. In addition, Huh et al. (2016) provide insights into the required quality of pre-training data. Our work is orthogonal to these studies, since we focus on studying weight generation, normalization, and aggregation techniques applied later on for new task adaptation.
@@
Although we investigate transfer learning scenarios, we review the imprinting applications and results from CL. Rebuffi et al. (2017) dynamically select a subset of examples for each class and update internal representations via gradient descent. They use a nearest mean classifier (NMC) with respect to the saved examples. Janson et al. (2022) use an NMC classifier as well and achieve good performance on CL benchmarks without any fine-tuning of the embeddings. However, they do not investigate the effect of normalization and using multiple proxies.
Findings of Prabhu et al. (2023) show that a simple, approximate $m$-nearest neighbor classifier outperforms existing methods in an Online CL setting when all data can be stored. In our work, however, we compare imprinting all data to a limited number of more representative proxies striving for efficiency.
@@
The phenomenon of NC was identified by Papyan et al. (2020) and refers to the convergence of the last-layer weight vectors to class means. It was shown that, regardless of the loss function, optimizer, batch-normalization, or regularization, NC will eventually occur (provided the training data has a balanced distribution) (Zhu et al., 2021; Han et al., 2022; Kothapalli, 2023). Nevertheless, complete neural collapse is practically unrealistic (Tirer et al., 2023). In transfer learning, Galanti et al. (2022) show that NC occurs on new samples and classes from the same distribution as the pre-training dataset, highlighting the usability of foundational models in such scenarios. In our work, we expand the survey on NC by experimenting with out-of-distribution classes belonging to different datasets and linking their degree of collapse to the success of certain imprinting strategies.
@@ 3.1 IMPRINT
We analyze the effect of weight generation (GEN), normalizations ($\texttt{NORM}=\{\texttt{NORM}\textsubscript{{pre}},\texttt{NORM}\textsubscript{{post}},\texttt{NORM}\textsubscript{{inf}}\}$), and aggregation (AGG). The IMPRINT framework depicted in fig.˜1 consists of three main building blocks: a foundation model FM, a weight generator GEN, and extendable classifier weights that are imprinted. The FM remains frozen throughout the experiments. It receives data from T as inputs and produces embedding vectors. The training process generates weight vectors for each of the $C$ classes in T. Hereby, embeddings from the FM are normalized before the generation (GEN) step according to NORMpre. The generated weight vectors per class are referred to as proxies, prototypes, or representatives (Movshovitz-Attias et al., 2017; Snell et al., 2017; Yang et al., 2018; Yu et al., 2020). These proxies are normalized according to NORMpost. As in the work by Qi et al. (2018), we do not use bias values. To classify the test data in T during inference, it is first embedded by the FM, normalized according to NORMinf, and finally aggregated by AGG, resulting in a predicted class label.
@@
The purpose of GEN is to determine how the embeddings of the training data in T are used to form the new weights. In contrast to Qi et al. (2018) which only incorporates one proxy per class (the mean), we add flexibility by allowing each class to have multiple proxies as in Mensink et al. (2013) and enable non-linear classification. We denote the number of proxies per class as $k$, ranging between $1$ and the number of samples, and investigate the following operations conducted per class to generate those:
@@
  • k-means: $k$-means cluster centers using KMeans from sklearn (Pedregosa et al., 2011). $k=1$ is the same as mean.
@@
quantile normalization (Amaratunga and Cabrera, 2001; Bolstad et al., 2003) can only be applied to generated weights. This non-linear operation distributes weights equally. Recall that if more than one class is contained in T ($c>1$), GEN is performed for each class consecutively, and the reference distribution changes accordingly. In particular, for the first class there is no reference distribution to map to. This is different from Hosoda et al. (2024), where new weights are matched to the distribution of the original classifier weights of the FM. Since we do not consider the classes used for pre-training the FM and especially do not assume access to their last-layer weights, this is not possible in our scenario.
@@ 3.2 Quantifying Neural Collapse
Neural collapse (NC) (Papyan et al., 2020) characterizes the state of the features produced by a classification neural network after training to near zero training loss. Namely, the learned embeddings of each class converge, i.e., collapse, to their class means. These globally centered class means and classifier weights form a simplex equiangular tight frame (ETF) – a collection of equal length and maximally equiangular vectors, that maximize the between-class variability. This results in an optimal linearly separable state for classification. In fig.˜3 (left), we illustrate the collapse of a FM on its pre-training data. The newly arrived data T from a different dataset is distributed more unevenly across the embedding space (right).
@@
Two important characteristics of NC are variability collapse, i.e., the within-class variability of the penultimate-layer embeddings collapses to zero, and convergence to nearest-mean-classification. We focus on variability collapse ($\mathcal{NC}_{1}$) as in Zhu et al. (2021). Given a foundation model FM and a finite dataset containing $C$ classes with (for simplicity) $N$ samples per class, we use its $l$-dimensional embeddings $\{\boldsymbol{\mathbf{h}}_{c,i}\}_{1\leq c\leq C,\;1\leq i\leq N}$ given by FM to define the global feature mean $\boldsymbol{\mathbf{h}}_{G}\coloneqq\frac{1}{CN}\sum_{c=1}^{C}\sum_{i=1}^{N}\boldsymbol{\mathbf{h}}_{c,i}$, class means $\overline{\boldsymbol{\mathbf{h}}}_{c}\coloneqq\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\mathbf{h}}_{c,i}$, the within-class covariance matrix $\mathbf{\Sigma}_{W}\coloneqq\frac{1}{CN}\sum_{c=1}^{C}\sum_{i=1}^{N}\bigl(\boldsymbol{\mathbf{h}}_{c,i}-\overline{\boldsymbol{\mathbf{h}}}_{c}\bigr)\bigl(\boldsymbol{\mathbf{h}}_{c,i}-\overline{\boldsymbol{\mathbf{h}}}_{c}\bigr)^{\top}$, and the between-class covariance matrix $\mathbf{\Sigma}_{B}\coloneqq\frac{1}{C}\sum_{c=1}^{C}\bigl(\overline{\boldsymbol{\mathbf{h}}}_{c}-\boldsymbol{\mathbf{h}}_{G}\bigr)\bigl(\overline{\boldsymbol{\mathbf{h}}}_{c}-\boldsymbol{\mathbf{h}}_{G}\bigr)^{\top}$ to finally compute
@@
Based on eq.˜1, an $\mathcal{NC}_{1}$ score closer to zero signifies a higher collapse. In contrast, an increase in multi-modality of data leads to a higher $\mathcal{NC}_{1}$ score (as analyzed in fig.˜11(a)). Note that this measurement is not independent of the embedding dimension $l$ and the number of classes $C$. According to NC, imprinting the mean, as originally done by Qi et al. (2018), is best when $\mathcal{NC}_{1}$ is small. We claim that when the data is not fully collapsed (as is often the case in practice), the scale of $\mathcal{NC}_{1}$ could guide the proxy generation method, e.g., having multiple proxies $k>1$ per class. We investigate this in section˜5.3.
@@ 3.3 Significance Testing with Critical Difference Diagrams
Our experiments compare large numbers of imprinting configurations across tasks with substantially different achievable accuracy levels. To obtain scale-independent statistical comparisons, we base our analysis on ranking (dis-)agreements rather than raw accuracy. Non-parametric rank-based tests are well suited for this setting. Following Demšar (2006), we use the Friedman test and Wilcoxon signed-rank tests because they do not assume normality or homoscedasticity, and can be applied uniformly to any evaluation metric. Empirical evidence in Demšar (2006) further suggests that these tests have good power in classifier comparison scenarios.
Concretely, for each evaluation instance (i.e., foundation model FM and task T) we compute the classification accuracy of all configurations and assign ranks (rank 1 = highest accuracy). This yields a matrix of ranks with one row per configuration and one column per evaluation instance. On this matrix we first apply the Friedman test to assess whether there are any overall differences between configurations. If the Friedman test rejects the global null hypothesis at significance level $\alpha=0.05$, we proceed with a post-hoc, pairwise analysis. Specifically, for every pair of configurations we run a two-sided Wilcoxon signed-rank test across evaluation instances. The resulting $p$-values are corrected for multiple comparisons using the Holm step-down procedure (Holm, 1979).